A ball is thrown downward from the top of a 46.0 m tower with an initial speed of 10.0 m/s. Assuming negligible air resistance, what is the speed of the ball just before hitting the ground?
1
Expert's answer
2020-09-16T10:09:20-0400

The kinematic law of the motion with the constant acceleration is:


"d = v_0t + \dfrac{gt^2}{2}"

where "d = 46m" is the height of the tower, "v_0 = 10m\/s" is the initial speed of the ball and "g = 9.81 m\/s^2" is the gravitational acceleration.

Solving the quadratic equation


"\dfrac{gt^2}{2} +v_0t - d = 0"

for "t", obtain the total time of falling:


"t_{fall} = \dfrac{-v_0 + \sqrt{v_0^2 + 2gd}}{2}"

The speed after "t_{fall}" seconds will be:


"v = v_0 + gt_{fall} = v_0 + \dfrac{g(\sqrt{v_0^2 + 2gd}-v_0)}{2} =\\
=10+ \dfrac{9.81(\sqrt{10^2 + 2\cdot 9.81\cdot 46}-10)}{2} \approx 116.26 m/s"

Answer. 116.26 m/s.